# ArchesClimate: Probabilistic Decadal Ensemble Generation With Flow Matching

### v1

@@ Abstract
Climate projections have uncertainties related to components of the climate system and their interactions. A typical approach to quantifying these uncertainties is to use climate models to create ensembles of repeated simulations under different initial conditions. Due to the complexity of these simulations, generating such ensembles of projections is computationally expensive. In this work, we present ArchesClimate, a deep learning-based climate model emulator that aims to reduce this cost. ArchesClimate is trained on decadal hindcasts of the IPSL-CM6A-LR climate model at a spatial resolution of approximately 2.5x1.25 degrees. We train a flow matching model following ArchesWeatherGen \@BBOPcitep\@BAP\@BBN(Couairon et al., 2024)\@BBCP, which we adapt to predict near-term climate. Once trained, the model generates states at a one-month lead time and can be used to auto-regressively emulate climate model simulations of any length. We show that for up to 10 years, these generations are stable and physically consistent. We also show that for several important climate variables, ArchesClimate generates simulations that are interchangeable with the IPSL model. This work suggests that climate model emulators could significantly reduce the cost of climate model simulations.
[…]
Journal of Advances in Modeling Earth Systems (JAMES)
ARCHES, INRIA, Paris, France UMR LOCEAN, IPSL, Sorbonne Université, IRD, CNRS, MNHN, Paris, France Department of Computer Science, University of Colorado Boulder, USA
[…]
Graham Clynegraham.clyne@inria.fr
{keypoints}
We present ArchesClimate, an AI model for probabilistically emulating a climate model at a monthly timescale
ArchesClimate simultaneously generates both oceanic and atmospheric dynamics that can be autoregressively generated for 10 years
We propose a set of evaluations of both the statistics and the physical properties of the AI-generated ensembles
Plain Language Summary
Climate modeling enables us to understand the impacts of a changing climate. Simulations from climate models can address a wide range of questions, including near-term (one to ten years) climate evolution to allow for informed and immediate policy decisions. To increase confidence in predictions from climate models, many simulations are integrated to form a probabilistic picture of the climate. These simulations require significant computational resources because of the complexity of the climate models used. Here we present ArchesClimate, a model to replicate the behavior of a climate model. Using machine learning, we are able to significantly reduce the computational cost of climate models. ArchesClimate generates both atmospheric and oceanic dynamics at a monthly temporal resolution and approximately 2.5x1.25 degree spatial resolution for up to 10 years.
@@ 1 Introduction
Ensemble generation is an important tool to investigate the climate at any timescale. Given the chaotic nature of the atmospheric and ocean dynamics, small variations in initial conditions can evolve to vastly different states. As such, ensembles of climate model simulations can be seen as distributions of possible future climates. By repeatedly sampling the simulated climate, we enable probabilistic analyses that support the investigation of a broad range of problems, including extreme event attribution and uncertainty quantification \@BBOPcitep\@BAP\@BBN(Fyfe et al., 2017; Fischer et al., 2023)\@BBCP.
@@
Furthermore, generating ensembles can significantly contribute to separating the internal variability from the effect of other sources of uncertainty, such as those associated with external forcings \@BBOPcitep\@BAP\@BBN(Maher et al., 2021; Eade et al., 2014)\@BBCP. Internal variability in climate models refers to the natural fluctuations in the climate system that occur without variations from external forcings (e.g. greenhouse gases or solar irradiation). These fluctuations arise from complex interactions between the atmosphere, ocean, land, and ice (e.g. El Niño or the North Atlantic Oscillation) — and can cause variations in climate at all time scales. Understanding internal variability improves confidence in detecting and projecting human-caused climate change.
Traditionally, ensemble members are generated by perturbing a set of initial conditions and running several instances of a climate model \@BBOPcitep\@BAP\@BBN(Deser et al., 2024)\@BBCP. This process is computationally expensive and thereby limits the widespread adoption of large ensemble generation \@BBOPcitep\@BAP\@BBN(Smith et al., 2019)\@BBCP. To reduce the cost of generating such ensembles, many atmospheric emulators have been made with increasing success. In \@BBOPcite\@BAP\@BBNBrenowitz et al. (2025)\@BBCP, the authors introduce a diffusion-based machine learning model trained on multiple atmospheric datasets to learn both weather and climate dynamics. While their approach demonstrates potential in generating realistic instantaneous states, it does not produce temporally consistent climatic sequences since the model is not auto-regressive, a seemingly critical requirement for long-term climate modeling. In contrast, \@BBOPcite\@BAP\@BBNWatt-Meyer et al. (2024)\@BBCP develops the ACE2 model, a deterministic emulator of ERA5 which can stably generate sequential states for up to 1000 years, emulating atmospheric dynamics that respond to forcings. While ACE2 marks a breakthrough in long-term atmospheric emulation, it is limited by its focus on the atmosphere alone, leaving coupled ocean–atmosphere processes and variability unresolved. Moreover, its deterministic design may underrepresent internal variability and uncertainty. Generalization beyond the present climate regime, particularly under future or paleoclimate conditions, also remains uncertain.
Several efforts have also been made to emulate oceanic components of climate models. \@BBOPcite\@BAP\@BBNGuo et al. (2024)\@BBCP proposes a global ocean emulator trained on climate simulations that produces stable decadal simulations. Complementing this approach, \@BBOPcite\@BAP\@BBNDheeshjith et al. (2025)\@BBCP introduces Samudra, a deterministic model that focuses on learning spatiotemporal coherence in ocean circulation patterns. While it performs well at depth and captures inter-annual variability, it does not respond to a changing climate.
@@
Each of these approaches incorporates design decisions tailored to their specific objectives, but our goal is distinct: to efficiently emulate the IPSL-CM6A-LR at decadal timescales using probabilistic modeling across both oceanic and atmospheric domains. We introduce ArchesClimate, an AI-driven probabilistic emulator that aims to learn climatic processes at a monthly temporal resolution. We adapted ArchesWeatherGen and trained it to emulate the IPSL-CM6A-LR model at approximately 2.5x1.25 degree resolution. We build on ArchesWeatherGen \@BBOPcitep\@BAP\@BBN(Couairon et al., 2024)\@BBCP, a state-of-the-art AI-based numerical weather prediction model, which is based on PanguWeather \@BBOPcitep\@BAP\@BBN(Bi et al., 2022)\@BBCP. ArchesWeatherGen provides a computationally efficient solution for ensemble generation that is adapted to climatic timescales.
@@
We use a dataset from the Decadal Climate Prediction Project (DCPP), a Model-Intercomparison Project aimed at improving predictability and understanding climate at the decadal timescale \@BBOPcitep\@BAP\@BBN(Boer et al., 2016)\@BBCP. It comprises ensembles of 10 members with a duration of 10 years, starting every year between 1960 and 2015. We use flow matching as our training scheme, a recent generative technique that learns a function to map one distribution to another distribution, usually from a multivariate normal distribution to the target data distribution \@BBOPcitep\@BAP\@BBN(Lipman et al., 2023)\@BBCP.
@@
We distinguish between an emulator of the input/output functionality of the climate model versus an emulator of how a climate model evolves in climate state. The former describes a machine learning model that takes in boundary conditions or forcings and outputs full climate states, replacing the functionality of a climate model. The latter describes a machine learning model that learns from the output of a climate model and learns dynamics from this output, often augmenting the emulated climate model. In this research, we do the latter, and therefore our setting uses data from the IPSL-CM6A-LR for both initial conditions and training data. If we worked directly from the forcings and did not use an initial state generated from IPSL-6CMA-LR, we could instead attempt to replace the functioning of IPSL-6CMA-LR.
@@
The manuscript is structured as follows: Section 2.1 and Section 2.2 present the dataset and architecture used in the research, Section 2.3 and Section 2.4 outline how we train and generate data, and Section 3 explains the experiments and their results. Section 4 discusses conclusions and next steps.
@@ 2.1 Dataset
We use generated states from the IPSL-CM6A-LR coupled climate model \@BBOPcitep\@BAP\@BBN(Boucher et al., 2020)\@BBCP submission to the DCPP \@BBOPcitep\@BAP\@BBN(Boer et al., 2016)\@BBCP. DCPP aims at exploring decadal climate prediction, its predictability and variability. We use data from hindcastA, an experiment that performs 10-year hindcasts starting from 1960 to the present. Ensembles are initialized every year on January 1st from 1960-2015. For example, the dataset contains a 10-member ensemble from 1960-1970 and another 10-member ensemble for 1961-1971. There is therefore significant temporal overlap between each ensemble. We use monthly outputs that are averaged values, corresponding to a dataset of approximately 70,000 simulated months. The 10 members have initial conditions generated by an assimilated run using observed sea surface temperature and sea surface salinity nudging over this period \@BBOPcitep\@BAP\@BBN(Estella-Perez et al., 2020; Servonnat et al., 2015; Deser et al., 2024)\@BBCP. This perturbation is enough to generate variability in the ensemble while keeping some features leading to decadal predictability unchanged. We will refer to the dataset as IPSL-DCPP hereafter.
@@
The variables laid out in Table 1 capture important atmospheric and oceanic dynamics, including the heat and water flux between the two domains. The variables net_flux (total heat exchange between atmosphere and ocean, see A for a detailed description) and evspsbl (evaporation) represent the interactions between the two domains. The resolution of the atmospheric model is 144x143 (lon x lat, roughly 2.5x1.25 degrees), and the resolution of the ocean model is 362x332 (lon x lat, roughly 1 degree). The oceanic variables are re-gridded onto the regular atmospheric grid using a nearest-neighbour interpolation. For oceanic variables, we mask grid points over land with zeros. We also use the atmospheric concentration of four greenhouse gases (CO2, CH4, CFC11eq, N2O) and solar irradiance. from input4mips, a data repository of standardized boundary conditions used in model-intercomparison projects \@BBOPcitep\@BAP\@BBN(Meinshausen & Nicholls, 2018; Lurton et al., 2020)\@BBCP.
Figure 1: On the right, a visualization of one state ($X_{t}$) from ISPL-DCPP, with surface and oceanic variables separated from atmospheric variables. External forcings are shown as a vector to the right.
@@ 2.2 Architecture
Here we describe the architecture of ArchesClimate. The backbone is identical to ArchesWeatherGen \@BBOPcitep\@BAP\@BBN(Couairon et al., 2024)\@BBCP. ArchesWeatherGen is a machine learning model based on PanguWeather that uses a hierarchical vision transformer and uses shifted-window attention for efficient and scalable modeling of visual features \@BBOPcitep\@BAP\@BBN(Bi et al., 2022; Liu et al., 2021; Vaswani et al., 2017)\@BBCP. ArchesWeatherGen uses an ensemble of deterministic models to predict the ensemble mean at the following timestep (6 hours) and generates a probabilistic ensemble of forecasts of the residual to predict the residual between the mean state and true state at the following timestep.
@@
  • We include greenhouse gas and solar irradiance forcings using conditional layer normalization as done in \@BBOPcite\@BAP\@BBNChen et al. (2021)\@BBCP. The forcings listed in Table 1 are included as parameters.
  • We remove axial attention to facilitate modeling of multiple domains. Axial attention breaks down attention into multiple 1D attentions along different axes (e.g., height and width separately for images). At the cost of increased complexity, we leave it to the model to determine the relationship between each domain and its interactions \@BBOPcitep\@BAP\@BBN(Ho et al., 2019)\@BBCP.
@@
  • We remove the skip connection to accommodate this large embedded dimension. While skip connections would preserve features and expressive capacity in the model, the total size of the model would exceed our computational resources.
  • We omit per-variable weighting in the loss function while keeping latitude weighting. Previous studies of weather emulators have included per-variable weighting \@BBOPcitep\@BAP\@BBN(Bi et al., 2022)\@BBCP, but given the different nature of our variable set, which contains ocean and atmospheric variables at a monthly timescale, we cannot reuse those weightings.
There are also changes to the training of ArchesClimate, which can be found in Section 2.3.
2.2.1 Forcings Included In ArchesClimate
To incorporate forcings in ArchesClimate, we adopt conditional layer normalization following \@BBOPcitep\@BAP\@BBN(Chen et al., 2021)\@BBCP. In this approach, each scalar forcing value is first passed through an embedding layer, which produces parameters that rescale and shift the outputs of a standard layer normalization. The forcings are CO2, CH4, CFC11eq, N2O and SSI as described in Table 1. We apply conditional layer normalization in all transformer blocks of ArchesClimate, enabling the model to integrate conditioning signals at both global and local levels. By introducing greenhouse gas forcings through this mechanism, the model can learn flexible relationships that adapt dynamically to different forcing scenarios.
@@ 2.3 Training
Following ArchesWeatherGen, we train both a deterministic and generative model to predict the next state at timestep $t+\delta$, where t is a timestep in IPSL-DCPP and $\delta$ is one month. A deterministic model, $f_{\theta}$, is trained to predict $X_{t+\delta}$ from $X_{t}$ where $X$ is a climatic state of IPSL-DCPP (see Figure 1). In this climate state, we include the forcings listed in Table 1 shown as $\text{Forcings}_{t}$ in Figure 2. They are included in both $f_{\theta}$ and $g_{\theta}$ via conditional layer normalization \@BBOPcitep\@BAP\@BBN(Chen et al., 2021)\@BBCP. We then train a generative model $g_{\theta}$ to predict the residual of the next state
$r_{t+\delta}=\frac{X_{t+\delta}-f_{\theta}(X_{t})}{\sigma}$ (1)
where $\sigma$ is the standard deviation of the residual ($X_{t+\delta}-f_{\theta}(X_{t})$) of the training dataset.
We use flow matching (FM) to train $g_{\theta}$. FM takes known distribution p and finds a path of probabilities to an unknown distribution q \@BBOPcitep\@BAP\@BBN(Lipman et al., 2023)\@BBCP. This probability path is discretized over $S\in\mathbb{N}$ steps, where S denotes the total number of discrete time intervals used to approximate the continuous flow from p to q. In our case, p is the Gaussian distribution and q is the distribution of the residual of the IPSL-DCPP. Training involves learning $\theta$ such that $g_{\theta}$ predicts $r_{t+\delta}$. The inputs to $g_{\theta}$ are the predicted state of the deterministic model $f_{\theta}(X_{t})$, the previous state $X_{t-\delta}$ and a residual noised according to a randomly chosen FM timestep $s$, $(1-s)r_{t+\delta}+s\epsilon$, where $\epsilon$ is noise sampled from a Gaussian distribution. During training, we sample s from a standard normal distribution, and use the sigmoid function as done in \@BBOPcite\@BAP\@BBNEsser et al. (2024)\@BBCP. See Figure 2 for a visualization of the process.
At each FM timestep $s$, the probability path is defined by a vector field that assigns a direction and magnitude to each point in our data to move between distributions. $g_{\theta}$ is updated to represent a vector field with the following loss function:
$\mathcal{L}=\mathbb{E}_{s\in\mathcal{U}(0,1),\epsilon\in\mathcal{N}(0,1)}\|(g_{\theta}(X_{t},f_{\theta}(X_{t}),X_{t-\delta},(1-s)r_{t+\delta}+s\epsilon)-(r_{t+\delta}-\epsilon)\|_{2}^{2}$ (2)
To recreate $X_{t+\delta}$ at a particular FM step $s$ we combine the output of the deterministic and generative model:
$X_{t+\delta,s+1}=f_{\theta}(X_{t}|X_{t-\delta})+g_{\theta}(X_{t},f_{\theta}(X_{t}),X_{t-\delta},\frac{(1-s)r_{t+\delta}+s\epsilon)}{\sigma})$ (3)
. We refer to \@BBOPcitep\@BAP\@BBN(Lipman et al., 2023; Esser et al., 2024)\@BBCP for more details on the flow matching process. There are two differences in the training between ArchesClimate and ArchesWeatherGen:
@@
  • The FM target is a vector field instead of a full sample (see Section 2.3 for more details). Predicting the full sample instead of the vector field caused instability in generated states, and we therefore predict the vector field when doing FM.
IPSL-DCPP has been split as follows: every ten years, the 10-member ensembles initialized in the years 1989 and 1999 and 2009 are held out as validation, and ensembles initialized in the years 1969,1979,2010-2015 are held out as the test set. We chose this test set to best target the task of interpolation. See C for a comparison of different train/test splits. We train the deterministic model for 10 hours on 4 A100 GPUs, and we train the probabilistic model for 20 hours on 4 A100 GPUs.

### v2

@@ Abstract
Climate projections have uncertainties related to components of the climate system and their interactions. A typical approach to quantifying these uncertainties is to use climate models to create ensembles of repeated simulations under different initial conditions. Due to the complexity of these simulations, generating such ensembles of projections is computationally expensive. In this work, we present ArchesClimate, a deep learning-based climate model emulator that aims to reduce this cost. ArchesClimate is trained on decadal hindcasts of the IPSL-CM6A-LR climate model at a spatial resolution of approximately 2.5x1.25 degrees. We train a flow matching model following ArchesWeatherGen (Couairon et al., 2024), which we adapt to predict near-term climate. Once trained, the model generates states at a one-month lead time and can be used to auto-regressively emulate climate model simulations. We show that for up to 10 years, these generations are stable and physically consistent. We also show that for several important climate variables, ArchesClimate generates simulations that are interchangeable with the IPSL model. This work suggests that climate model emulators could significantly reduce the cost of climate model simulations.
@@ 1 Introduction
Ensemble generation is an important tool to investigate the climate at any timescale. Given the chaotic nature of the atmospheric and ocean dynamics, small variations in initial conditions can evolve to vastly different states. As such, ensembles of climate model simulations can be seen as distributions of possible future climates. By repeatedly sampling the simulated climate, we enable probabilistic analyses that support the investigation of a broad range of problems, including extreme event attribution and uncertainty quantification (Fyfe et al., 2017; Fischer et al., 2023).
@@
Furthermore, generating ensembles can significantly contribute to separating the internal variability from the effect of other sources of uncertainty, such as those associated with external forcings (Maher et al., 2021; Eade et al., 2014). Internal variability in climate models refers to the natural fluctuations in the climate system that occur without variations from external forcings (e.g. greenhouse gases or solar irradiation). These fluctuations arise from complex interactions between the atmosphere, ocean, land, and ice (e.g. El Niño or the North Atlantic Oscillation) — and can cause variations in climate at all time scales. Understanding internal variability improves confidence in detecting and projecting human-caused climate change.
Generating climate model ensembles traditionally requires perturbing initial conditions and running multiple resource-intensive instances of complex climate models (Deser et al., 2024). This practice is computationally expensive and severely limits the widespread generation of very large ensembles (Smith et al., 2019). To overcome this hurdle, significant research has focused on developing efficient atmospheric and coupled system emulators that can drastically reduce the associated computational cost.
Significant progress has been made in probabilistic climate emulation, notably with models like NeuralGCM (Kochkov et al., 2024). This model operates as a hybrid, physics-based, deep learning auto-regressive emulator, successfully generating temporally consistent atmospheric sequences. NeuralGCM captures robust variability by introducing the necessary randomness through the injection of noise directly into the neural network component of its hybrid architecture. Another approach is Spherical DYffusion (Cachay et al., 2023), a diffusion-based model explicitly designed as a probabilistic climate model emulator. This contrasts with the model cBottle, introduced by Brenowitz et al. (2025), a generative diffusion model that successfully learns the short-term, instantaneous dynamics of weather and climate when trained on multiple atmospheric datasets, demonstrating the potential for generating realistic atmospheric states; however, its non-autoregressive design means it does not produce the long-term, temporally consistent sequences necessary for extended climate modeling. A major advance in auto-regressive climate modeling is seen in ACE and its successor, ACE2 (Watt-Meyer et al. (2024, 2023)). ACE2 is a deterministic emulator capable of generating stable sequential atmospheric states for up to 1000 years. However, its deterministic design may inherently underrepresent crucial internal variability, and its exclusive focus on the atmosphere omits the essential dynamics of the coupled ocean–atmosphere system. Additionally, the LUCIE framework (Guan et al., 2024) provides an approach aimed at efficiently learning and predicting components of the uncoupled system using ERA5 data. Addressing the need for a coupled approach, several models have emerged. DLESyM (Cresswell-Clay et al., 2025) is a complex coupled atmosphere-ocean emulator demonstrating the practical feasibility of simultaneously modeling the two major components trained with ERA5 data. The SamudrACE model (Duncan et al., 2025) comlements this by combining the deterministic ocean model Samudra (Dheeshjith et al., 2025) with ACE2, resulting in a model that performs well and shows effective exchange between the ocean and atmosphere. Similarly, ACE-SOM2 (Clark et al., 2024) couples ACE2 with a simpler single-layer ocean model to investigate model response under different forcings, addressing generalization beyond stable climate scenarios.
@@
Each of these approaches incorporates design decisions tailored to their specific objectives, but our goal is distinct: to efficiently emulate the IPSL-CM6A-LR at decadal timescales using probabilistic modeling across both oceanic and atmospheric domains in order to generate realistic ensemble members. We introduce ArchesClimate, an AI-driven probabilistic emulator that aims to learn climatic processes at a monthly temporal resolution. We adapted ArchesWeatherGen and trained it to emulate the IPSL-CM6A-LR model at approximately 2.5x1.25 degree resolution. We build on ArchesWeatherGen (Couairon et al., 2024), a state-of-the-art AI-based numerical weather prediction model, which is based on PanguWeather (Bi et al., 2022). ArchesWeatherGen provides a computationally efficient solution for ensemble generation that is adapted to climatic timescales.
@@
We use a dataset from the Decadal Climate Prediction Project (DCPP), a Model-Intercomparison Project aimed at improving predictability and understanding climate at the decadal timescale (Boer et al., 2016). It comprises ensembles of 10 members with a duration of 10 years, starting every year between 1960 and 2015. This dataset provides us with many samples per year, which allows us to efficiently learn the dynamics of IPSL-CM6A-LR over decadal scales. We use flow matching as our training scheme, a recent generative technique that learns a function to map one distribution to another distribution, usually from a multivariate normal distribution to the target data distribution (Lipman et al., 2023).
We target a monthly resolution to enhance the computational efficiency of the emulation process. This decision draws upon the theoretical framework that long-term system dynamics are driven by large-scale, low-frequency variability (Hasselmann, 1976; Held et al., 2010). Therefore, we prioritize the representation of these slower modes by using a monthly timestep.
@@
We distinguish between an emulator of the input/output functionality of the climate model versus an emulator of how a climate model evolves in climate state. The former describes a machine learning model that takes in boundary conditions or forcings and outputs full climate states, replacing the functionality of a climate model. The latter describes a machine learning model that learns dynamics from the output of a climate model, often augmenting the emulated climate model. In this research, we do the latter, and therefore our setting uses data from the IPSL-CM6A-LR for both initial conditions and training data. If we worked directly from the forcings and did not use an initial state generated from IPSL-6CMA-LR, we could instead attempt to replace the functioning of IPSL-6CMA-LR.
@@
The manuscript is structured as follows: in section 2, subsection 2.1 and subsection 2.2 present the dataset and architecture used in the research, subsection 2.4 and subsection 2.5 outline how we train and generate data, and section 3 explains the experiments and their results. section 4 discusses conclusions and next steps.
@@ 2.1 Dataset
We use generated states from the IPSL-CM6A-LR coupled climate model (Boucher et al., 2020) submission to the DCPP (Boer et al., 2016). The DCPP aims at exploring decadal climate prediction, its predictability and variability. We use data from hindcastA, an experiment that performs 10-year hindcasts starting from 1960 to the present. Ensembles are initialized every year on January 1st from 1960-2015. For example, the dataset contains a 10-member ensemble from 1960-1970 and another 10-member ensemble for 1961-1971. While both these runs contain the years 1961-1970, they will have their own internal variability and exhibit different dynamics. Therefore for any given initialization year, the 10-member ensemble resulting from this year will be different from the years produced by other initializations. This is an important element of this dataset, allowing us to learn the dynamics of the IPSL-CM6A-LR over a smaller timeframe with significantly more samples than in other MIP-style experiments. We use monthly outputs that are averaged values, corresponding to a dataset of approximately 70,000 simulated months. The 10 ensemble members have initial conditions generated by an assimilated run using observed sea surface temperature and sea surface salinity nudging over this period (Estella-Perez et al., 2020; Servonnat et al., 2015; Deser et al., 2024). Each ensemble member is forced with perturbed sea surface temperature and sea surface salinity. This perturbation is enough to generate variability in the ensemble while keeping some features leading to decadal predictability unchanged. We will refer to the dataset as IPSL-DCPP hereafter.
@@
The variables laid out in Table 1 capture important atmospheric and oceanic dynamics, including the heat and water flux between the two domains. The variables net_flux (total heat exchange between atmosphere and ocean, see A for a detailed description) and evspsbl (evaporation) represent the interactions between the two domains. The resolution of the atmospheric model is 144x143 (lon x lat, roughly 2.5x1.25 degrees), and the resolution of the ocean model is 362x332 (lon x lat, roughly 1 degree). The oceanic variables are re-gridded onto the regular atmospheric grid using a nearest-neighbour interpolation. For oceanic variables, we mask grid points over land with zeros. We also use the normalized, globally averaged atmospheric concentration of four greenhouse gases (CO2, CH4, CFC11eq, N2O) and solar irradiance from input4mips, a data repository of standardized boundary conditions used in model-intercomparison projects (Meinshausen and Nicholls, 2018; Lurton et al., 2020).
Figure 1: On the right, a visualization of one state ($X_{t}$) from ISPL-DCPP, with surface and oceanic variables separated from atmospheric variables. Globally averaged and normalized external forcings are shown as a vector to the right.
@@ 2.2 Architecture
Here we describe the architecture of ArchesClimate. The backbone is very similar to ArchesWeatherGen, with the differences listed below (Couairon et al., 2024). The ArchesWeather architecture consists of a two-part machine learning system for weather forecasting: a deterministic model and a generative model. The deterministic component is built on a 3D Swin U-Net transformer structure based on PanguWeather and employs a novel Cross-Level Attention (CLA) mechanism to efficiently handle interactions across different vertical pressure levels (Bi et al., 2022; Liu et al., 2021; Vaswani et al., 2017). The second component, ArchesWeatherGen, is a probabilistic generative model that shares the same architecture and uses a flow matching approach to model the residual weather states, thereby enabling the generation of reliable ensemble forecasts to quantify uncertainty. This combined approach is designed to enhance both the accuracy of the deterministic prediction and the robustness of probabilistic forecasting while being computationally efficient. ArchesWeatherGen operates in a two-stage process to generate its forecasts. First, it uses an ensemble of deterministic models to calculate the most likely future atmospheric conditions, or the ensemble mean, for the next 6-hour timestep. Second, instead of generating a probabilistic forecast of the entire next state, the model focuses its computational power on generating a probabilistic ensemble of residuals. The residual represents the difference between the predicted ensemble mean and the actual state. By focusing on predicting this residual, the model efficiently captures uncertainty and internal variability around the most likely forecast.
@@
  • We include greenhouse gas and solar irradiance forcings using conditional layer normalization as done in Chen et al. (2021). The forcings listed in Table 1 are included as parameters.
  • We remove axial attention to facilitate modeling of multiple domains. Axial attention breaks down attention into multiple 1D attentions along different axes (e.g., height and width separately for images). At the cost of increased complexity, we remove axial attention and attend to the flattened atmospheric layers, leaving it to the model to determine the relationship between each domain and its interactions (Ho et al., 2019).
@@
  • We omit per-variable weighting in the loss function while keeping latitude weighting. Previous studies of weather emulators have included per-variable weighting (Bi et al., 2022), but given the different nature of our variable set, which contains ocean and atmospheric variables at a monthly timescale, we cannot reuse those weightings.
There are also changes to the training of ArchesClimate, which can be found in subsection 2.4.
2.3 Forcings Included In ArchesClimate
To incorporate forcings in ArchesClimate, we adopt conditional layer normalization following (Chen et al., 2021). In this approach, each scalar forcing value is first passed through an embedding layer, which produces parameters that rescale and shift the outputs of a standard layer normalization. The forcings are CO2, CH4, CFC11eq, N2O and SSI as described in Table 1. We apply conditional layer normalization in all transformer blocks of ArchesClimate, enabling the model to integrate conditioning signals at both global and local levels. By introducing greenhouse gas forcings through this mechanism, the model may be able to learn flexible relationships that adapt dynamically to different forcing scenarios. This idea is explored in Appendix B but extensive investigation is left for further work.
@@ 2.4 Training
Following ArchesWeatherGen, we train both a deterministic and generative model to predict the next state at timestep $t+\delta$, where t is a timestep in IPSL-DCPP and $\delta$ is one month. A deterministic model, $f_{\theta}$, is trained to predict $X_{t+\delta}$ from $X_{t}$ where $X$ is a climatic state of IPSL-DCPP (see 1). In this climate state, we include the forcings listed in Table 1 shown as $\text{Forcings}_{t}$ in Figure 2. They are included in both $f_{\theta}$ and $g_{\theta}$ via conditional layer normalization (Chen et al., 2021). We then train a generative model $g_{\theta}$ to predict the residual of the predicted next state:
$r_{t+\delta}=\frac{X_{t+\delta}-f_{\theta}(X_{t+\delta}|X_{t},X_{t-\delta})}{\sigma}$ (1)
where $\sigma$ is the standard deviation of the residuals ($X_{t+\delta}-f_{\theta}(X_{t+\delta}|X_{t},X_{t-\delta})$) of the training dataset.
We use flow matching (FM) to train $g_{\theta}$. FM takes known distribution p and finds a path of probabilities to an unknown distribution q (Lipman et al., 2023). This probability path is discretized over $S\in\mathbb{N}$ steps, where S denotes the total number of discrete time intervals used to approximate the continuous flow from p to q. In our case, p is the Gaussian distribution and q is the distribution of the residual of the IPSL-DCPP. Training involves learning $\theta$ such that $g_{\theta}$ predicts $r_{t+\delta}$. The inputs to $g_{\theta}$ are the predicted state of the deterministic model $f_{\theta}(X_{t})$, the previous state $X_{t-\delta}$ and a residual noised according to a randomly chosen FM timestep $s\in S$, $(1-s)r_{t+\delta}+s\epsilon$, where $\epsilon$ is noise sampled from a Gaussian distribution. During training, we sample s from a standard normal distribution $s\sim sigmoid(\mathcal{N}(0,1))$, and use the sigmoid function as done in Esser et al. (2024). See Figure 2 for a visualization of the process.
At each FM timestep $s$, the probability path is defined by a vector field that assigns a direction and magnitude to each point in our data to move between distributions. $g_{\theta}$ is updated to represent a vector field of the residual $(r_{t+\delta}-\epsilon)$ with the following loss function:
$\mathcal{L}=\mathbb{E}_{s\in\mathcal{U}(0,1),\epsilon\in\mathcal{N}(0,1)}\|(g_{\theta}(X_{t},f_{\theta}(X_{t+\delta}|X_{t},X_{t-\delta}),\frac{(1-s)r_{t+\delta}+s\epsilon}{\sigma})-(r_{t+\delta}-\epsilon)|_{2}^{2}$ (2)
We refer to (Lipman et al., 2023; Esser et al., 2024) for more details on the flow matching process.
@@
The IPSL-DCPP dataset, which consists of 55 10-year, 10-member ensembles (one initialized annually), was carefully split to enable a direct comparison of full ensemble predictions with ArchesClimate. The split holds out 3 ensembles (initialized in 1989, 1999, and 2009) as the validation set and 8 ensembles (initialized in 1969, 1979, and 2010–2015) as the test set, leaving the remaining 44 ensembles for training. This configuration is justified by the ensembles’ distinct initial conditions, despite any temporal overlap, which allows for robust evaluation by comparing a complete 10-member ensemble from a given initialization year against the output of ArchesClimate. See C for a comparison of different train/test splits. We train the deterministic model for 10 hours on 4 A100 GPUs, and we train the probabilistic model for 20 hours on 4 A100 GPUs. See subsection 2.6 for a list of hyperparameters used.